OpenStudy (anonymous):

Just checkin my answer: Express the radical in simplest form: the square root of 1/ 11 A. square root of 11 B. square root of 11/11 C. 1/square root of 11

OpenStudy (anonymous):

I think it's B

OpenStudy (istim):

B is wrong...

OpenStudy (anonymous):

 this is what B looks like

OpenStudy (unklerhaukus):

\[\sqrt{1\over 11} =\sqrt{1\over 11} \times \sqrt{\frac{{11}} {11}} \]\[= \frac {\sqrt{11}}{11}\] \[B.\]

OpenStudy (anonymous):

I think you're right

OpenStudy (anonymous):

Yea i think i'm right too...Istim is wrong

OpenStudy (unklerhaukus):

\[B. =C.\] i think C. is a simpler

OpenStudy (anonymous):

@UnkleRhaukus It's usually improper to have a radical in the denominator though. sqrt(1/11) = sqrt(1)/sqrt(11) = 1/sqrt(11) Multiply by sqrt(11)/sqrt(11) which is the same as 1. sqrt(11)/[ sqrt(11) * sqrt(11) ] Squareroots cancel out. sqrt(11)/11 Answer is B.

OpenStudy (unklerhaukus):

well yeah

OpenStudy (unklerhaukus):

\[=11^{-1/2}\] looks even simpeler to me